OpenStudy (anonymous):

Does the rule 4x4^x represent a linear or an exponential function? Explain.

OpenStudy (anonymous):

Exponential. x raised to any power higher than 1 is exponential.

OpenStudy (campbell_st):

the rule for the same base is add the powers \[x^1\times x^4 =x^5\] your problem uses the same rule \[4^1\times 4^x =4^{x+1}\]

OpenStudy (campbell_st):

so it is an exponential function